find the vertex, the line of symmetry, the max or min value of the quadratic function
f(x)=x^2-2x-3

Answer by nerdybill(7384)

find the vertex, the line of symmetry, the max or min value of the quadratic function
f(x)=x^2-2x-3
.
"axis of symmetry":
x = -b/(2a)
x = -(-2)/(2*1)
x = 2/(2)
x = 1 ("line of symmetry")
.
Plug the above in to find the vertex:
f(x)=x^2-2x-3
f(1)=1^2-2(1)-3
f(1)=1-2-3
f(1)=-4
.
Because the coefficient associated with the x^2 term is POSITIVE, the vertex gives the MINIMUM:
Therefore, minimum value of quadratic is -4